OpenStudy (bekkah323):

-4^2+6 I got 22 but the book said the answer is -10?

jimthompson5910 (jim_thompson5910):

-4^2 is not 16 because you square 4 first, then you make it negative

jimthompson5910 (jim_thompson5910):

you are thinking of (-4)^2

jimthompson5910 (jim_thompson5910):

so it should be this -4^2+6 -16+6 -10

OpenStudy (bekkah323):

oh, ok thank you!

jimthompson5910 (jim_thompson5910):

the best way to think of it is to think of -4^2 as -1*4^2 PEMDAS says to do exponents first, then multiplication, so -1*4^2 = -1*16 = -16 if you use parenthesis to change the order of operations, you get (-1*4)^2 = (-4)^2 = (-4)*(-4) = 16

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (bekkah323):

got it

jimthompson5910 (jim_thompson5910):

ok great